OpenStudy (anonymous):

Translate the following phrases into algebraic expressions. Explain, in complete sentences, the difference between the two. • twice the sum of x and y • the sum of twice x and y

OpenStudy (anonymous):

1.. x plus (add) y squared (multiply by itself) x + y*y = x + y^2 2. 8 times(multiply) a number, x, decreased (subtract) by y 8x - y 3. the sum (add) of x and y, squared x + y*y =x + y^2 4. 3 times(multiply) the difference (subtract) of x and y =3( x + y) 5. 6 less than (subtract) x cubed 6 - x*x*x =6 - x^3 * twice the sum(addition) of x and y =2(x + y) * the sum(addition) of twice x and y =2x + y First means two times x added to 2 times y Second means two times x added to y only. There is a difference of a y between them.